# Controllable Image Colorization with Instance-aware Texts and Masks

### v1

@@ Abstract
Index Terms:
Image colorization, diffusion models, instance level, segmentation masks, textual descriptions, dataset.
Figure 1: Precise and flexible instance-aware colorization. MT-Color can respect: a) generate pleasing unconditional colorization results automatically, b) colorize grayscale images in an instance-aware manner with user-provided instance masks and instance texts. The generation process of MT-Color preserves pixel information and achieve strong color-text binding.
I Introduction
@@
II Related Work
II-A Automatic Colorization
@@
II-B Text-based Colorization
@@
II-C Diffusion-based Colorization
@@ III Methodology
III-A Pixel-level mask attention mechanism
@@
III-B Instance mask and text guidance module
@@
Given a latent representation $f_{x}\in\mathbb{R}^{h\times w\times c}$, a set of instance masks $M_{n}=\{m_{k}\}_{k=1}^{n}\in\{0,1\}^{n\times H\times W}$ and a set of instance texts $T_{n}=\{\tau_{k}\}_{k=1}^{n}\in\mathbb{R}^{n\times l_{t}}$, where $l_{t}$ represents the maximum possible length of each instance text, we first convert the instance masks and texts into instance representations that can be input to the instance mask and text guidance module. For instance texts, we use the pre-trained CLIP text encoder to transform $T_{n}$. For instance masks, we use a multi-layer perceptron (MLP) to extract their features. The MLP consists of 3 convolutional layers. We concatenate the corresponding feature of each instance and pass it through a MLP composed of three fully connected layers. As 3 shows, this process yields the instance feature set $\Gamma_{n}=\{\gamma_{k}\}_{k=1}^{n}\in\mathbb{R}^{n\times l_{\gamma}}$, where $l_{\gamma}$is the length of the instance representation.
@@
[…]
to obtain a self-attention map, denoted as $self\_map$, where $Q_{p}=W_{Q_{p}}\cdot p$,$K_{p}=W_{K_{p}}\cdot p$,$V_{p}=W_{V_{p}}\cdot p$. The size of $self\_map$ is $(l_{x}+n)\times(l_{x}+n)$.
@@
$M_{\text{self}}(i,j)=\begin{cases}m_{k},&\exists k\in\{1,\ldots,n\},\\ &m_{k}(w_{i},h_{i})=m_{k}(w_{j},h_{j})=1\\ 0,&\text{otherwise}\end{cases}$ (5)
@@
$\hat{f_{x}}=(M_{sel\mathit{f}\_map}\circ self\_map\cdot V_{p})[1:l_{x},1:l_{x}]$ (7)
Here, we only take the results from the latent feature part, specifically the portion $[1:l_{x},1:l_{x}]$, as the output of the instance mask and text guidance module.
III-C Multi-instance sampling for inference
@@
III-D Dataset construction pipeline
@@
IV Experiments
(o) Input
(p) CT2
(q) DDColor
(r) DeOldify
(s) L-CAD
(t) Ours
(u) GT
@@ IV Experiments
IV-A Model Training
IV-A1 Training Strategy
@@
$\mathcal{L}_{2}=\|\epsilon-\epsilon_{\Theta}(z_{t},t,\tau_{g},c,M_{n},T_{n})\|^{2}_{2}$ (10)
@@ IV-A2 Training Settings
To improve model robustness and support both conditional and unconditional colorization, we randomly set the input mask and text to null tokens with a probability of 50%. For multi-instance sampling, we set $\alpha=0.2$ and $\beta=0.2$.
All training is performed on 4 NVIDIA A40 GPUs. The first stage is trained for 25,000 iterations, followed by 20,000 iterations in the second stage.
IV-B Comparison with prior work
@@
IV-B1 Quantitative Comparison
TABLE I: Quantitative comparison for unconditional colorization on GPT-color. $\uparrow$($\downarrow$)indicates higher(lower) is better. Best performances are highlighted in bold and second best performances are highlighted in underline. Our model performs well on non-reference human perceptual-level metrics.
@@ IV-B1 Quantitative Comparison
We benchmark our method against previous methods on GPT-color and report quantitative results in Table I. It is worth noting that the metrics widely used in previous works like PSNR, SSIM and FID [42] mainly focus on the structural similarity between images. However, since images with a high structural similarity to the original image may not necessarily conform to the natural image distribution and human perception, we believe that using perceptual-level metrics is necessary in the task of image colorization. Thus, we introduce 4 perceptual-level non-reference image quality assessment (NR_IQA) metrics, NIQE [43], MANIQA [44], MUSIQ[45] and TOPIQ_NR [46] to assess our method. We found that our model did not achieve state-of-the-art performance on metrics that reflect the structural similarity since these metrics do not focus on whether the generated images are colorful or realistic. Moreover, MT-Color’s output is of higher resolution than other diffusion-based methods’ output, which leads to worse pixel-level metric results. In terms of Colorfulness [47], our model performs well, which indicates that our model is able to produce colorful results. On human perceptual-level metrics, our model performs better than other models, indicating that the colorful images generated by our model are more in line with natural distribution patterns and human visual perception.
IV-B2 Qualitative Comparison
The qualitative comparison results are shown in Figure 5. We observed that DeOldify, as a GAN-based model, suffers from large areas of muted colors and a lack of color variety, resulting in poor visual quality. CT2 and DDcolor, as Transformer-based colorization models, produce more vivid and varied colors but exhibit color bleeding issues. Additionally, these models often apply different colors to the same object, such as the sign in the second row, leading to unrealistic results. Both L-CAD and our model are diffusion-based, whose results exhibit almost no color bleeding, with overall vibrant and natural colors in the images. Our model provides a more diverse color palette, such as the colorful sugar needles on the bread in the first row. Moreover, the resolution of MT-Color’s results are fixed to $512\times 512$, which is clearer than the $256\times 256$ resolution of L-CAD and CT2.More results and analysis are shown in the supplementary material.
IV-B3 Comparison with Other Diffusion-Based Methods
TABLE II: Summary of advantages of MT-Color over existing diffusion-based methods.
@@ IV-B3 Comparison with Other Diffusion-Based Methods
Diffusing Colors[31] 256$\times$256 $\times$ $\times$ $\times$
Piggybacked[32] 256$\times$256 $\times$ $\times$ $\times$
L-CAD[25] 256$\times$256 $\times$
Ours 512$\times$512
Several recent works leverage the generative power of pre-trained diffusion models for image colorization. However, due to the lack of open-source implementations for many of these methods, we are unable to conduct direct qualitative and quantitative comparisons. Instead, Table II provides a summary comparison between these approaches and our proposed MT-Color.
@@
IV-C Ablation Study
IV-C1 Pixel-Level Masked Attention Mechanism
We conduct an ablation study to evaluate the effectiveness of the proposed Pixel-Level Masked Attention Mechanism (PMAM). The quantitative results are summarized in Table III, and visual comparisons are presented in Figure 6.
@@ IV-C1 Pixel-Level Masked Attention Mechanism
TABLE III: Quantitative results of ablation study for PMAM.
@@
IV-C2 Instance Mask and Text Guidance Module
@@
We further conduct quantitative evaluations by computing the CLIP-score [7] on the GPT-Color test set. Each instance is cropped using its mask, and the CLIP-score is calculated between the cropped region and its corresponding text. As shown in Table IV, the complete model achieves the highest score, indicating stronger alignment between generated colors and textual descriptions.
TABLE IV: Quantitative results of ablation study on instance mask and text guidance module.
@@
IV-C3 Multi-Instance Sampling Strategy
@@
Quantitative results in Table V show that the complete multi-instance sampling method achieves the highest CLIP-score and MUSIQ, indicating improved semantic alignment and perceptual quality. These results validate the necessity of separately sampling and fusing instance-level results with region-aware cropping.
TABLE V: Quantitative results of ablation study on the multi-instance sampling strategy.
@@
IV-C4 GPT-Color Comparison
TABLE VI: Comparison between GPT-Color and other datasets.
@@ IV-C4 GPT-Color Comparison
Currently, two mainstream COCO-based datasets are used for image colorization: COCO-Stuff and Multi-instance. COCO-Stuff is primarily designed for instance segmentation, where the global text annotations are manually written, and the instance-level annotations are limited to category labels, lacking detailed color information. Multi-instance is tailored for colorization tasks, where global text is generated by BLIP, but it does not provide instance-level textual descriptions. As summarized in Table VI, GPT-Color combines the strengths of both datasets—it supports automatic text generation, includes rich color information, and provides fine-grained instance-level descriptions.
TABLE VII: Quantitative comparison of textual descriptions among datasets using CLIP-Score.
@@
We demonstrate several visual samples for qualitatively comparison in the supplemental material. To quantitatively assess the quality of textual annotations, we compute the CLIP-Score between global text and images across the three datasets. For instance-level evaluation, we apply instance masks from COCO-Stuff and GPT-Color to extract individual instance regions and then compute the CLIP-Score between the cropped image patches and their corresponding instance descriptions. Results are shown in Table VII.
@@
TABLE VIII: Performance comparison of models trained on different datasets.
@@
We further evaluate the training capability of each dataset by training the same model on COCO-Stuff, Multi-instance, and GPT-Color, and testing on the GPT-Color test set. Since Multi-instance does not provide instance-level text, we supply empty instance texts during training for fair. The results are shown in Table VIII.

### v2

@@ Abstract
keywords:
[…]
Figure 1: MT-Color can respect: a) generate pleasing unconditional colorization results automatically, b) colorize grayscale images in an instance-aware manner with user-provided instance masks and instance texts. The generation process of MT-Color preserves pixel information and achieve strong color-text binding.
1 Introduction
@@
2 Related Work
2.1 Automatic colorization
@@
2.2 Text-based colorization
@@
2.3 Diffusion-based colorization
@@ 3 Methodology
In this section, we first introduce the MT-Color’s base method, diffusion models and ControlNet. We then propose the pixel-level masked attention mechanism, which is responsible for the pixel-level fusion of conditional grayscale image representations and latent representations. Next, we detail the instance mask and text guidance module, which integrates instance representations with corresponding tokens in latent features. Moreover, we introduce the multi-instance sampling strategy, which enhances the independence of each instance during the sampling process. Lastly we introduce the vision-language model-aided automatic construction pipeline of our instance-level image colorization dataset, GPT-color.
3.1 Preliminary
3.1.1 Diffusion models
Diffusion models consist of a forward noising process and a reverse denoising process. In the forward process, Gaussian noise $\epsilon$ is gradually added to the clean data sample $x_{0}$ over $T$ time steps, resulting in a sequence of progressively noised samples $x_{1},\dots,x_{T}$. The reverse process aims to recover $x_{0}$ from a noisy input $x_{t}$ by learning a denoising model $\epsilon_{\theta}$ that predicts the noise added at each time step $t$.
To reduce the computational cost of diffusion models in pixel space, the Latent Diffusion Model (LDM) performs the diffusion process in a compressed latent space. Given an optional condition $c$, the training objective of LDM is defined as:
$\mathcal{L}_{\text{LDM}}=\mathbb{E}_{x_{0},\epsilon,t,c}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t},t,c)\right\|_{2}^{2}\right],$ (1)
where $z_{t}$ denotes the latent representation at time step $t$, and $\epsilon_{\theta}$ is the denoising network.
3.1.2 ControlNet
ControlNet is a neural network architecture designed to introduce explicit conditional control into pretrained text-to-image diffusion models. It constructs a deep and expressive encoder by creating a trainable copy of selected layers from the base LDM. This copy learns to encode additional control signals, while the original model remains mostly fixed.
The trainable branch and the original model are linked through "zero convolution" layers, which help suppress the propagation of harmful noise during training. The training objective of ControlNet-augmented LDM can be formulated as:
$\mathcal{L}_{cond}=\mathbb{E}_{x_{0},\epsilon,t,c,y}\left[\left\|\epsilon-\epsilon_{\theta}(z_{t},t,c,y)\right\|_{2}^{2}\right],$ (2)
where $c$ denotes the textual condition, and $y$ represents the additional structural condition provided by ControlNet.
3.2 Pixel-level masked attention mechanism
@@
3.3 Instance mask and text guidance module
@@
Given a latent representation $f_{x}\in\mathbb{R}^{h\times w\times c}$, a set of instance masks $M_{n}=\{m_{k}\}_{k=1}^{n}\in\{0,1\}^{n\times H\times W}$ and a set of instance texts $T_{n}=\{\tau_{k}\}_{k=1}^{n}\in\mathbb{R}^{n\times l_{t}}$, where $l_{t}$ represents the maximum possible length of each instance text, we first convert the instance masks and texts into instance representations that can be input to the instance mask and text guidance module. For instance texts, we use the pre-trained CLIP text encoder to transform $T_{n}$. For instance masks, we use a multi-layer perceptron (MLP) to extract their features. The MLP consists of 3 convolutional layers. We concatenate the corresponding feature of each instance and pass it through a MLP composed of three fully connected layers. As 5 shows, this process yields the instance feature set $\Gamma_{n}=\{\gamma_{k}\}_{k=1}^{n}\in\mathbb{R}^{n\times l_{\gamma}}$, where $l_{\gamma}$is the length of the instance representation.
@@
$\text{self\_map}=\text{Softmax}(\frac{Q_{p}K_{p}^{T}}{\sqrt{d}})$ (6)
to obtain a self-attention map, denoted as self_mapself\_map, where $Q_{p}=W_{Q_{p}}\cdot p$$K_{p}=W_{K_{p}}\cdot p$$V_{p}=W_{V_{p}}\cdot p$. The size of self_mapself\_map is $(l_{x}+n)\times(l_{x}+n)$.
@@
$M_{\text{self}}(i,j)=\begin{cases}m_{k},&\exists k\in\{1,\ldots,n\},m_{k}(w_{i},h_{i})=m_{k}(w_{j},h_{j})=1\\ 0,&\text{otherwise}\end{cases}$ (7)
@@
$\hat{f_{x}}=(M_{\text{self\_map}}\circ\text{self\_map}\cdot V_{p})[1:l_{x},1:l_{x}]$ (9)
Here, we only take the results from the latent feature part, specifically the portion $1:l_{x},1:l_{x}]$, as the output of the instance mask and text guidance module.
3.4 Multi-instance sampling for inference
@@
3.5 Dataset construction pipeline
@@
4 Experiments
@@
4.1 Training strategy
@@
$\mathcal{L}=\mathbb{E}_{x_{0},\epsilon,t,c,\tau_{g},M_{n},T_{n}}\|\epsilon-\epsilon_{\Theta}(z_{t},t,\tau_{g},c,M_{n},T_{n})\|^{2}_{2}$ (12)
@@ 4.2 Experiment settings
To improve model robustness and support both conditional and unconditional colorization, we randomly set the input mask and text to null tokens with a probability of 50%. All training is performed on 4 NVIDIA A40 GPUs. The first stage is trained for 25,000 iterations, followed by 20,000 iterations in the second stage.
4.3 Comparison with prior work
@@
4.3.1 Quantitative comparison
Table 1: Quantitative comparison for unconditional colorization on GPT-color. $\uparrow$($\downarrow$)indicates higher(lower) is better. Best performances are highlighted in bold and second best performances are highlighted in underline. Our model performs well on non-reference human perceptual-level metrics.
@@ 4.3.1 Quantitative comparison
We benchmark our method against previous methods on GPT-color and report quantitative results in Table 1. It is worth noting that the metrics widely used in previous works like PSNR, SSIM and FID [8] mainly focus on the structural similarity between images. However, since images with a high structural similarity to the original image may not necessarily conform to the natural image distribution and human perception, we believe that using perceptual-level metrics is necessary in the task of image colorization. Thus, we introduce 4 perceptual-level non-reference image quality assessment (NR_IQA) metrics, NIQE [23], MANIQA [39], MUSIQ[14] and TOPIQ_NR [4] to assess our method. We found that our model did not achieve state-of-the-art performance on metrics that reflect the structural similarity since these metrics do not focus on whether the generated images are colorful or realistic. Moreover, MT-Color’s output is of higher resolution than other diffusion-based methods’ output, which leads to worse pixel-level metric results. In terms of Colorfulness [9], our model performs well, which indicates that our model is able to produce colorful results. On human perceptual-level metrics, our model performs better than other models, indicating that the colorful images generated by our model are more in line with natural distribution patterns and human visual perception.
4.3.2 Qualitative comparison
The qualitative comparison results are shown in Figure 5. We observed that DeOldify, as a GAN-based model, suffers from large areas of muted colors and a lack of color variety, resulting in poor visual quality. CT2 and DDcolor, as Transformer-based colorization models, produce more vivid and varied colors but exhibit color bleeding issues. Additionally, these models often apply different colors to the same object, such as the sign in the second row, leading to unrealistic results. Both L-CAD and our model are diffusion-based, whose results exhibit almost no color bleeding, with overall vibrant and natural colors in the images. Our model provides a more diverse color palette, such as the colorful sugar needles on the bread in the first row. Moreover, the resolution of MT-Color’s results are fixed to $512\times 512$, which is clearer than the $256\times 256$ resolution of L-CAD and CT2.More results and analysis are shown in the appendix.
4.3.3 Comparison with other diffusion-based methods
Table 2: Summary of advantages of MT-Color over existing diffusion-based methods.
@@ 4.3.3 Comparison with other diffusion-based methods
Diffusing Colors[42] $\times256$ $\times$ $\times$ $\times$
Piggybacked[21] $\times256$ $\times$ $\times$ $\times$
L-CAD[36] $\times256$ $\times$
Ours $\times512$
Several recent works leverage the generative power of pre-trained diffusion models for image colorization. However, due to the lack of open-source implementations for many of these methods, we are unable to conduct direct qualitative and quantitative comparisons. Instead, Table 2 provides a summary comparison between these approaches and our proposed MT-Color.
@@
4.4 Ablation study
4.4.1 Pixel-level masked attention mechanism
We conduct an ablation study to evaluate the effectiveness of the proposed Pixel-Level Masked Attention Mechanism (PMAM). The quantitative results are summarized in Table 3, and visual comparisons are presented in Figure 6.
@@ 4.4.1 Pixel-level masked attention mechanism
Table 3: Quantitative results of ablation study for PMAM.
@@
4.4.2 Instance mask and text guidance module
@@
We further conduct quantitative evaluations by computing the CLIP-score [26] on the GPT-Color test set. Each instance is cropped using its mask, and the CLIP-score is calculated between the cropped region and its corresponding text. As shown in Table 4, the complete model achieves the highest score, indicating stronger alignment between generated colors and textual descriptions.
Table 4: Quantitative results of ablation study on instance mask and text guidance module.
@@
4.4.3 Multi-instance sampling strategy
@@
Quantitative results in Table 5 show that the complete multi-instance sampling method achieves the highest CLIP-score and MUSIQ, indicating improved semantic alignment and perceptual quality. These results validate the necessity of separately sampling and fusing instance-level results with region-aware cropping.
Table 5: Quantitative results of ablation study on the multi-instance sampling strategy.
@@
4.4.4 Dataset comparison
Table 6: Comparison between GPT-Color and other datasets.
@@ 4.4.4 Dataset comparison
Currently, two mainstream COCO-based datasets are used for image colorization: COCO-Stuff and Multi-instance. COCO-Stuff is primarily designed for instance segmentation, where the global text annotations are manually written, and the instance-level annotations are limited to category labels, lacking detailed color information. Multi-instance is tailored for colorization tasks, where global text is generated by BLIP, but it does not provide instance-level textual descriptions. As summarized in Table 6, GPT-Color combines the strengths of both datasets—it supports automatic text generation, includes rich color information, and provides fine-grained instance-level descriptions.
Table 7: Quantitative comparison of textual descriptions among datasets using CLIP-Score.
@@
We demonstrate several visual samples for qualitatively comparison in the appendix. To quantitatively assess the quality of textual annotations, we compute the CLIP-Score between global text and images across the three datasets. For instance-level evaluation, we apply instance masks from COCO-Stuff and GPT-Color to extract individual instance regions and then compute the CLIP-Score between the cropped image patches and their corresponding instance descriptions. Results are shown in Table 7.
@@
Table 8: Performance comparison of models trained on different datasets.
@@
We further evaluate the training capability of each dataset by training the same model on COCO-Stuff, Multi-instance, and GPT-Color, and testing on the GPT-Color test set. Since Multi-instance does not provide instance-level text, we supply empty instance texts during training for fair. The results are shown in Table 8.
@@ 5 Conclusion
6 Acknowledgment
This work is supported by National Natural Science Foundation of China (62271308), STCSM(24ZR1432000, 24511106902, 24511106900, 22DZ2229005), 111 plan (BP0719010), and State Key Laboratory of UHD Video and Audio Production and Presentation.
Appendix A Discussion on hyperparameters
To explore the effect of the two key hyperparameters $\alpha$ and $\beta$, we conduct a series of experiments and report the results in Table 9. Here, $\alpha$ controls the portion of individual sampling steps, while $\beta$ adjusts the weight of global noise during sampling.
Table 9: Hyperparameters comparison study.
[…]
0 0 0.2162 22.36 11.41
0.2 0 0.2240 21.07 15.87
0.2 0.2 0.2273 23.12 11.39
0.2 0.4 0.2124 22.76 13.78
0.4 0.4 0.2305 20.81 17.76
From the results in Table 9, we observe that both $\alpha$ and $\beta$ play an important role in balancing semantic alignment, reconstruction fidelity, and realism. We find that higher values of $\alpha$ and lower $\beta$ enhance the binding between instances and texts, but often lead to unstable or inconsistent generation quality. When both hyperparameters are disabled ($\alpha=0,\beta=0$), the model yields moderate performance across all three metrics, serving as a baseline without multi-instance sampling.
Setting both $\alpha$ and $\beta$ to 0.4 results in the highest CLIP-Score (0.2305), but the lowest PSNR and the worst FID (17.76), indicating a significant trade-off: although text-image alignment improves, the generated images become less faithful and perceptually coherent.
When both $\alpha$ and $\beta$ are set to 0.2, the model achieves the best balance: the highest PSNR (23.12) and a competitive FID (11.39), while maintaining a strong CLIP-Score (0.2273).We therefore choose this setting as our default during inference.
Appendix B Visual dataset comparison
To further demonstrate the advantages of GPT-Color in generating textual descriptions, we randomly selected several images from the COCO dataset and compared the corresponding global text descriptions from GPT-Color, COCO-Stuff, and Multi-Instance datasets, as shown in Figure 9.
We observe that the global text in COCO-Stuff is notably brief, focusing primarily on the general scene and covering only a limited subset of the objects present in the image. Moreover, it lacks detailed color information, which is essential for image colorization tasks. In contrast, the Multi-Instance dataset provides longer descriptions that mention more objects than COCO-Stuff. However, the descriptions often contain irrelevant or non-informative sentences, such as rhetorical questions or repetitive mentions of the same object (e.g., “side of a floater,” “part of a floater,” and “edge of a boat” in the third image). Additionally, despite being tailored for image colorization, Multi-Instance does not consistently provide color details for every mentioned object.
In comparison, GPT-Color begins with a description of the overall color tone of the image, followed by a comprehensive enumeration of objects within the scene, each annotated with specific color information. This structure ensures both completeness and relevance in the text.
From a qualitative perspective, the textual annotations in GPT-Color are more informative, coherent, and better suited for guiding colorization tasks than those in COCO-Stuff and Multi-Instance.
Appendix C More visual results of conditional colorization
In this section, we present several examples from the GPT-Color validation set to further demonstrate the conditional colorization capabilities of our proposed MT-Color model, which leverages global text descriptions, instance segmentation masks, and instance-level textual annotations.
As shown in Figure 10, MT-Color is capable of producing precise and diverse instance-aware image colorization guided by user-provided global descriptions, instance masks, and instance texts. The colorization results not only adhere closely to the color constraints specified by the instance-level inputs, but also align well with human visual perception. Furthermore, the output resolution is increased to $512\times 512$, offering clearer and more visually appealing results.
Benefiting from the multi-instance sampling strategy, the color of each object is influenced not only by the corresponding instance text, but also by the global description. For instance, in the case where the towel held by a girl is not annotated with an instance mask or description, its color is still correctly inferred based on the global text input.
However, due to the inherent stochasticity of diffusion models, MT-Color may occasionally fail to preserve fine pixel-level details from the grayscale input or even generate suboptimal results. Addressing this limitation remains an open direction for future work.
Figure 9: Qualitative comparison of global textual descriptions across GPT-Color, COCO-Stuff, and Multi-Instance datasets.
Figure 10: Visual examples of conditional colorization with global texts, instance masks and instance texts on GPT-color.
Appendix D More comparisons of automatic colorization
Table 10: Quantitative comparison for unconditional colorization on COCO-Stuff dataset. $\uparrow$($\downarrow$) indicates higher(lower) is better. Best performances are highlighted in bold and second best performances are highlighted in underline.
COCO-Stuff Pixel-level Perceptual-level
[…]
Deoldify 25.7857 23.2442 0.8677 15.3988 3.6135 69.9131 0.5004 0.5961
DDColor 35.4759 22.9509 0.8614 9.7483 3.5588 69.5143 0.4910 0.5981
L-CAD 28.8897 24.2191 0.8728 11.3773 4.9759 54.5308 0.4021 0.6283
CT2 40.4601 23.0503 0.8692 12.5434 4.7118 56.3082 0.4361 0.5278
Ours 33.0894 23.0729 0.8704 11.9967 4.1312 70.4190 0.5113 0.6358
Table 11: Quantitative comparison for unconditional colorization on ImageNet dataset. $\uparrow$($\downarrow$) indicates higher(lower) is better. Best performances are highlighted in bold and second best performances are highlighted in underline
ImageNet Pixel-level Perceptual-level
[…]
Deoldify 25.6872 23.6722 0.9107 7.8766 4.5813 68.7172 0.5314 0.6512
DDColor 40.7589 22.6318 0.8911 5.136 4.5918 68.7083 0.5258 0.6520
L-CAD 25.2565 22.4030 0.8690 11.0019 5.4117 58.6934 0.4512 0.6497
CT2 40.1252 23.2567 0.8760 11.8491 5.3477 60.4682 0.4769 0.6006
Ours 35.3771 22.7644 0.8779 10.7835 3.9466 69.2562 0.5255 0.6556
As shown in Figure 11 and 12, we provide more unconditional colorization visual results of our model and the comparison with previous methods on COCO-Stuff dataset and ImageNet dataset, respectively. Meanwhile, we test our model and previous methods on these two datasets and report quantitative results in Table 10 and 11.
Since the generated images of CT2 are cropped and resized, in quantitative experiments we crop and resize the ground truth images to match the generated images. As is shown, the results of DeOldify suffer from dull tones and uninspiring colors. Although CT2 and L-CAD could generate colorful and visual appealing images, the resolution of their outputs is limited to $256\times 256$, which is too low for human visual perception. Since DDColor is based on Transformer architecture and is proposed solely for automatic colorization, it could generate good colorization results while preserving details of the grayscale images and the original resolution. However, DDColor sometimes generate uneven colors due to lack of semantic information. Our proposed MT-Color could not only generate natural and colorful images that better match human visual perception, but also preserve some semantic information learned from the specific-designed dataset and fix the resolution to $512\times 512$, which is clearer for human perception. Nonetheless, MT-Color sometimes could not preserve pixel details due to the stochasticity of diffusion models.
Figure 11: Qualitative comparison results for unconditional colorization. All examples are from COCO-Stuff.
Figure 12: Qualitative comparison results for unconditional colorization. All examples are from ImageNet-5k.
Appendix E Discussion on computation costs
We extend the original ControlNet architecture from latent space to pixel space and adopt a multi-instance sampling strategy during inference to enable precise instance-aware colorization. While these improvements increase computational costs, they significantly enhance performance. Table 12 presents the computational details of MT-Color and a comparison with L-CAD, a diffusion-based baseline. MT-Color is trained on an NVIDIA A40 GPU, while L-CAD is trained on an NVIDIA RTX 3090 GPU. All inference is conducted using the NVIDIA A40 GPU.
The introduction of a pixel-level attention mechanism and a more sophisticated guidance module significantly increases the number of parameters in MT-Color, nearly doubling those of L-CAD. During inference, MT-Color produces images with a resolution of $512\times 512$, which contains four times as many pixels as L-CAD’s $256\times 256$ output. Consequently, the inference time of MT-Color is approximately four times that of L-CAD due to the additional computational cost introduced by pixel-level attention and multi-instance sampling. Nonetheless, this cost is acceptable given the improvements in output resolution and instance-level precision. MT-Color’s memory usage during inference is around 18.0 GB, which remains within the range of high-end consumer GPUs.
Table 12: Computation details of MT-Color and L-CAD under $\alpha=0.1$.
Method Parameters Training Inference
Time Memory Time per Image Memory Resolution
L-CAD 1052M 120h 14.3GB 15.2s 7.3GB $256\times 256$
MT-Color (Ours) 1950M 124h 40.7GB 68.6s 18.0GB $512\times 512$
[…]